# Super-Resolution with Structured Motion

### v1

@@ 6 Experimental Results
We implemented the deconvolution algorithms in python using numpy for the Wiener deconvolution method and both cvxpy and pylops for computing reconstructions using sparsity priors. We found that using the OSQP solver within cvxpy yields good results with small images, as seen in the simulations throughout the paper. For the experiments with real data, we used the split-bregman algorithm implemented in pylops due to the computational demands of processing larger images.
@@ 6.1 Setup
The data used in the following reconstructions consists of sets of images captured by a monochrome 1.6 megapixel (1440x1080) CMOS camera (Thorlabs Zelux) with 3.45 $\mu$m square pixels and a wide-angle 8 mm f1.8 Edmund Optics lens opened to maximum aperture and focused at approximately 3 meters. This particular choice of lens and sensor leads to an optical system that is not diffraction limited despite the relatively small pixel size.
The camera was mounted on a movable stage and controlled by a nanopositioning system (Nanomotion II) that uses dc stepper motors to move the stage at nanometer level resolution in two separately controlled axes. Movements made by the stage have high precision and repeatability, though no position readout is available. The stage is connected to and controlled by a computer via a serial port.
@@ 6.2 Grid Exposures
In our setup we move the camera instead of the sensor. To acquire a static grid of images, the stage was moved through a series of positions in a plane parallel to the target, coming to a stop at each position and waiting for two frames to be captured, before continuing. The overall grid motion was repeated five times. Averaging many frames reduces the effects of sensor noise and running multiple trials of the grid sequence minimizes the effects of temporal illumination variation during the acquisition.
@@
With grid capture (see Section 6.2), the image of the target moves by less than $\Delta$ in both the vertical and horizontal directions, with pictures taken at $f$ sub-pixel positions, for a total of $f\times f$ image locations. Taking into account the magnification $m$, the distance the stage travels between neighboring grid positions, i.e. step size, is $\frac{\Delta}{fm}$.
For $f=8$, we use 64 grid locations. The magnification of our system was calculated to be 0.002484 (we use a wide-angle lens), which gives a step size of 0.17361 mm.
Due to the large field of view of the camera, our targets occupy a relatively small (150x190) area in the images acquired, which means that for the experiments presented here, we are only using a small region of the physical sensor in the camera. In other applications, the same setup would allow for capturing very wide field of views.
Figure 9 shows several results, each a small ROI of the target area. We note that in each example we can see high-resolution details appear in the super-resolution images (e.g. punctuation and to some extent serifs, elements in USAF group -1, the serrated top edge of the ketchup packet).
(a) Static image
@@
Figure 10 shows that our magnification model holds for three-dimensional objects that protrude out of the target plane. The front of the LEGO(R) bus is approximately 15 cm away from the calibrated target plane and high-resolution details on the minifigures are still resolved.
(a) Static image
@@
Figure 10: Super-resolution using an 8x8 grid of static images. The bottom row shows a crop of the front section of the bus taken from the full image shown in the top row. Using a moving camera instead of a moving sensor requires the camera motion to be proportional to the distance from the object. Nonetheless we can image three-dimensional objects that are relatively close to a target plane. In this case, a constant magnification model is still accurate.
@@ 6.3 Scan Motion
Since our stage is slow, we used a relatively long exposure to capture images with motion blur. The exposure was determined by how long it takes for the image on the sensor to move by $\Delta$ (the size of the sensor pixel). The maximum velocity of the stage is only 2.5 mm/s and therefore, for these experiments we used an exposure time of about five hundred milliseconds.
Due to the fact that the stage has no position readout capabilities and the camera cannot be synchronized with it, the initial camera position for each acquisition must be estimated. The printed targets include a black rectangle, which provides a step edge that is used for the position estimation; similar to the blurred point source shown in Figure 5, a sharp edge can be used for localization. The exposure time, vertical position, and velocity of the stage during each capture are known, so once the initial position has been estimated, the trajectory of motion during the acquisition is fully defined.
@@ 7 Conclusion and Future Work
We described several methods for increasing the limits of super-resolution with a moving sensor/camera.
Using a grid of static images, the ambiguity in super-resolution is restricted to relatively few Fourier coefficients. This enables fast reconstruction using a Wiener filter, and higher effective super-resolution factors with sparse priors and compressed sensing techniques.
@@
Note that by using two different magnifications during acquisition, or a sensor with multiple pixel sizes, one could potentially remove all of the ambiguity in the super-resolution problem. Figure 12 illustrates the Fourier transform of a discrete periodic box of width 4 and similarly of a box of width 7. We see that the zeros of the two Fourier transforms occur in different places; if we were to record the convolution of a high-resolution image with boxes of two sizes, we would eliminate essentially all ambiguity in the super-resolution problem. Taking pictures while moving at different speeds could also be used to integrate the brightness function over regions of different sizes.
We have also shown that motion blur can in fact be used improve the resolution with a single image. This leads to the idea that to take a sharper picture one should shake or vibrate the camera. In synthetic experiments, we have shown that simply vibrating the image sensor during capture can lead to higher resolution images. The key here is that we solve a motion deblurring problem at the same time that we increase the image resolution, which can be accomplished as long as we have high-resolution trajectory information.

### v2

@@ 6 Experimental Results
We implemented the reconstruction algorithms in python. We used numpy for the Wiener deconvolution method and both cvxpy and pylops for computing reconstructions using sparse image priors. We found that using the OSQP solver within cvxpy yields good results with small images, as seen in the simulations throughout the paper. For the experiments with real data, we used the split-bregman algorithm implemented in pylops due to the computational demands of processing larger images.
@@
All of the running times scale linearly with the image sizes. A direct reconstruction of a very large (gigapixel) image using a sparse image prior would be impractical with our current implementation. However, we could break the image into (overlapping) tiles and reconstruct each tile independently in parallel.
@@ 6.1 Setup
The data used in the following reconstructions consist of sets of images captured by a monochrome 1.6 megapixel (1440x1080) CMOS camera (Thorlabs Zelux) with 3.45 $\mu$m square pixels and a wide-angle 8 mm f1.8 Edmund Optics lens opened to maximum aperture and focused at approximately 3 meters. This particular choice of lens and sensor leads to an optical system that is not diffraction limited despite the relatively small pixel size.
The pixels in our camera are smaller than many professional photography cameras, which are also not diffraction limited when using large apertures. Pixel shift technology that is available on various cameras could in principle be used to implement the imaging systems described here, although that would likely require special access to the control system of the camera.
For the experiments described here, instead of moving the sensor, we move the camera. The camera was mounted on a movable stage and controlled by a nanopositioning system (Nanomotion II) that uses stepper motors to move the stage at nanometer level resolution in two separately controlled axes. Movements made by the stage have high precision and repeatability, though no position readout is available. The stage is connected to and controlled by a computer via a serial port.
@@
Our system is highly controlled via a stationary target, fixed lighting, and long acquisition times. One application where these conditions are available is in digital archiving. In this field, unique camera architectures have been developed to capture high-resolution images (see, e.g. [4]).
@@ 6.2 Grid Exposures
To acquire a static grid of images, the stage was moved through a series of positions in a plane parallel to the target, coming to a stop at each position and waiting for two frames to be captured, before continuing. The overall grid motion was repeated five times. Averaging many frames reduces the effects of sensor noise and running multiple trials of the grid sequence minimizes the effects of temporal illumination variation during the acquisition.
@@
With grid capture (see Section 6.2), the image of the target moves by $\frac{\Delta}{f}$ between adjacent grid locations, for a total of $f\times f$ locations. Taking into account the magnification $m$, the distance the stage travels between neighboring grid positions, i.e. step size, is $\frac{\Delta}{fm}$.
For $f=8$, we use 64 grid locations. The magnification of our system was calculated to be 0.002484 (we use a wide-angle lens), which gives a step size of 0.1736 mm.
Due to the large field of view of the camera, our targets occupy a relatively small (150x190) area in the images acquired, which means that, for the experiments presented here, we are only using a small region of the physical sensor in the camera. In other applications, the same setup would allow for capturing very wide field of views.
Figure 10 shows several results, each a small ROI of the target area. We note that in each example we can see high-resolution details appear in the super-resolution images, such as the punctuation and to some extent serifs in the text, elements in USAF group -1 in the resolution target, and the serrated top edge of the ketchup packet.
(a) Static image (1 of 64)
@@
Moving the camera instead of the sensor can lead to motion parallax. Figure 11 shows that a constant magnification model can still be used for three-dimensional objects that protrude out of the target plane. The front of the LEGO(R) bus is approximately 15 cm away from the calibrated target plane and high-resolution details on the minifigures are still resolved. Even with twice as much depth variation, the parallax effect would be small due to the large distance between the camera and the object.
Note that in Figure 11, the striped overalls of a LEGO minifigure appear uniform in a static image (a) but have visible stripes in the super-resolution image (c). The stripes can be seen accurately as lines with widths 3-4x smaller than the original pixel size, indicating a 3-4x improvement.
(a) Static image (1 of 64)
@@
Figure 11: Super-resolution using an 8x8 grid of static images. The bottom row shows a crop of the front section of the bus. Using a moving camera, instead of a moving sensor, requires the camera motion to be proportional to the distance from the object. Nonetheless, we can image three-dimensional objects that are relatively close to a target plane. In this case, a constant magnification model is still accurate.
For the examples above, we averaged multiple frames taken from each camera position to minimize the effect of measurement noise and to determine an upper limit on the reconstruction quality obtained using a physical system. Figure 12 compares a reconstruction obtained using a single frame from each position to a reconstruction obtained using ten frames from each position. While the reconstruction using a single frame from each position has more noise, it still recovers high-resolution features.
(a) 1 frame per position
(b) 10 frames per position
Figure 12: By averaging 10 frames per position, we reduce the effect of measurement noise (TV reconstructions).
@@ 6.3 Scan Motion
Since our stage is slow, we used a relatively long exposure to capture images with motion blur. The exposure was determined by how long it takes for the image on the sensor to move by $\Delta$ (the sensor pixel size). The maximum velocity of the stage is only 2.5 mm/s and therefore, for these experiments we used an exposure time of about five hundred milliseconds.
Due to the fact that the stage has no position readout capabilities and the camera cannot be synchronized with it, the initial camera position for each acquisition must be estimated. The exposure time, vertical position, and velocity of the stage during each capture are known. The printed targets include a black rectangle, which provides a step edge that is used for the horizontal position estimation; similar to the blurred point source shown in Figure 5, the initial position of a sharp edge can be localized to sub-pixel location from a moving camera.
@@ 7 Conclusion and Future Work
We described several results related to the limits of super-resolution with a moving sensor/camera.
Our contributions include:
1) A new analysis and demonstration of super-resolution and deconvolution with a box using sparse image priors.
2) The idea and demonstration that motion blur can be beneficial for super-resolution.
@@
We have shown that by using a grid of static images, the ambiguity in super-resolution can be restricted to relatively few Fourier coefficients. It is interesting to note that by using multiple magnifications, or a sensor with multiple pixel sizes, one could potentially remove all of the ambiguity in the super-resolution problem. Figure 14 illustrates the Fourier transform of a discrete periodic box of width 4 and similarly of a box of width 7. We see that the zeros of the two Fourier transforms occur in different places; if we were to record the convolution of a high-resolution image with boxes of two sizes, we would eliminate essentially all ambiguity in the super-resolution problem.
The idea that motion blur can be used to improve resolution is somewhat counterintuitive and leads to the notion that to take a sharper picture, one should shake or vibrate the camera. In synthetic experiments, we have shown that simply vibrating the image sensor during capture can lead to higher resolution images. In this case reconstruction involves solving a deconvolution problem while increasing resolution at the same time.